# VA-GS: Enhancing the Geometric Representation of Gaussian Splatting via View Alignment

### v1

@@ 1 Introduction
Existing methods have attempted to enhance the geometric capabilities of Gaussian splatting. For example, SuGaR [14] constructs a density field from Gaussians and extracts meshes via level-set searching, but it struggles with large smooth surfaces and is computationally expensive. 2DGS [16] models scenes using 2D oriented planar Gaussian disks, which inherently represent surfaces and provide view-consistent geometry. However, 2DGS has difficulty reconstructing background geometry and often produces incomplete or distorted surfaces in complex or unbounded scenes. GOF [56] constructs an opacity field from Gaussians and extracts surfaces using Marching Tetrahedra [10], yielding adaptive mesh resolution without volumetric fusion. Nonetheless, thin structures can be lost and strong lighting contrasts still cause artifacts. GS-Pull [60] integrates a neural signed distance field (SDF), dynamically pulling Gaussians toward the zero-level set of the learned SDF. While this improves surface completeness, it introduces additional network complexity, produces overly smooth surfaces, and primarily focuses on foreground object reconstruction. PGSR [4] fits Gaussians to local planar hypotheses and uses unbiased depth rendering to improve geometric accuracy. However, it does not fully resolve the challenges posed by complex lighting and remains sensitive to boundary ambiguities in non-planar regions. Overall, previous methods have introduced geometric regularizers or hybrid representations and achieved significant progress. However, they still struggle to address two persistent challenges: illumination-induced artifacts (e.g., shadows and specular highlights) and accurate surface boundary delineation, as shown in Fig. 1. Illumination effects distort photometric losses, while ambiguous boundaries often result in geometry drift or holes.
@@ 2 Related Work
View Synthesis and Gaussian Splatting. Neural Radiance Fields (NeRF) [31] pioneered high-fidelity novel view synthesis by representing a scene as a continuous volumetric density and view-dependent color field, optimized via differentiable volume rendering. Subsequent works accelerated training and rendering through hybrid representations such as multi-resolution hash grids [32], explicit voxel or sparse tensor grids [40, 2], and learned feature planes [53]. However, these volumetric methods still entail high memory and computational costs. 3DGS [21] departs from dense volumes by modeling a scene as a sparse cloud of anisotropic 3D Gaussians. Follow-up work has enhanced visual fidelity through anti-aliasing and level-of-detail control [55, 38], improved training speed and robustness under sparse views using density regularization and learned radiance priors [46, 34], and extended 3DGS to dynamic scenes [29, 47], relighting [13], and animation [51]. Geometry-aware variants such as FatesGS [17], DNGaussian [23] and GeoGaussian [26] address sparse-view and textureless regions, while methods like Instantsplat [11] and Scaffold-GS [28] accelerate convergence by leveraging pretraining or hybrid implicit-explicit designs. Despite these advances, most 3DGS variants primarily emphasize appearance quality and lack mechanisms to enforce explicit surface geometry, motivating dedicated reconstruction techniques.
@@
Surface Reconstruction with Gaussians. Extracting accurate surfaces from a 3DGS representation is challenging due to its unstructured nature and supervision based solely on RGB signals. Early approaches convert Gaussians into volumetric density or opacity fields: SuGaR [14] builds a density field and applies level-set search with Poisson reconstruction [20], but it struggles to recover large, smooth surfaces; GOF [56] accumulates per-view alpha values into an opacity volume and extracts iso-surfaces with Marching Tetrahedra [10], achieving adaptive resolution but often missing fine, thin structures under high lighting contrast. Other methods project Gaussians into oriented 2D disks (surfels) and fuse via Truncated Signed Distance Function (TSDF) fusion [33] or Poisson reconstruction. 2DGS [16] and GSurfels [8] improve local alignment but tend to introduce distortions in unbounded scenes and result in incomplete background geometry. PGSR [4] fits Gaussians to planar patches and adds multi-view photometric and geometric regularization, excelling on planar man-made scenes but remaining sensitive to non-planar boundaries. More recent works [54, 60, 30, 1, 24, 25] integrate Signed Distance Fields (SDF) to guide Gaussian placement. GSDF [54] and 3DGSR [30] jointly optimize a neural SDF branch alongside Gaussian parameters using volume-rendered depth and normal supervision, which improves surface smoothness but requires additional network branches. GS-Pull [60] leverages SDF gradients to pull Gaussians toward the zero-level set, enhancing alignment at the cost of limiting object-level reconstruction and producing overly smooth results. Methods that incorporate depth or normal estimators [5, 57, 41, 45, 43] impose priors on Gaussians but rely on TSDF fusion’s fixed resolution or Poisson reconstruction’s sensitivity to noisy inputs, and often struggle under varying illumination or around complex geometric boundaries. Our approach enforces view consistency through multi-faceted constraints during Gaussian optimization, enabling high-fidelity mesh extraction even under challenging lighting and boundary conditions.
@@ 3 Preliminaries
3DGS [21] explicitly represents a scene as a collection of anisotropic 3D Gaussians, which can be rendered to images from arbitrary viewpoints using a splatting-based rasterization technique [61]. Specifically, each 3D Gaussian $\mathbb{G}$ is defined as:
@@
Normal and Depth Estimation from Gaussians. The covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ of a 3D Gaussian can be decomposed into a rotation matrix $\bm{R}$ and a scaling matrix $\bm{S}$, i.e., $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$, where $\bm{R}$ contains the three orthogonal eigenvectors, and $\bm{S}$ encodes the scale along these directions. This decomposition resembles an ellipsoid representation: the eigenvectors define the axes of the ellipsoid, while the scale values correspond to the axis lengths. As optimization progresses, the initially spherical Gaussian flattens and approaches a plane [19]. We take the direction corresponding to the smallest scale factor as the normal $\bm{n}$ of the Gaussian. The distance from the local plane to the camera center is then computed as $\bm{d}=(\bm{R}_{c}^{\top}(\bm{\mu}-\bm{T}_{c}))^{\top}(\bm{R}_{c}^{\top}\bm{n})$, where $\bm{R}_{c}$ is the rotation from the camera to the world frame, and $\bm{T}_{c}$ is the camera center in world coordinates. Given the normal and distance, the depth is obtained by intersecting the viewing ray with the local plane: $\bm{z}=\bm{d}/(\bm{R}_{c}^{\top}\bm{n}~\bm{K}^{-1}\bar{\bm{p}})$, where $\bar{\bm{p}}$ is the homogeneous coordinate of the pixel (we use $\bm{p}$ to denote both the homogeneous and 2D pixel coordinates for simplicity), and $\bm{K}$ is the intrinsic matrix of the camera. Finally, the per-pixel distance, depth, and normal maps under the current viewpoint are rendered using $\alpha$-blending as defined in Eq. (2), where the attribute color $\bm{c}_{i}$ is replaced with the corresponding Gaussian attributes.
@@ 4.1 Single-View Alignment
Edge-aware Image Reconstruction. The original 3DGS [21] and its variants typically employ a color rendering loss, which combines the L1 reconstruction error with a D-SSIM term. While effective for overall image quality, this loss alone is insufficient for accurately capturing object boundaries during surface reconstruction, and it tends to overly smooth high-frequency regions and complex structures. To address this limitation, we propose an edge-aware image reconstruction loss that encourages the model to better preserve sharp structures and boundary details:
@@
Normal-based Geometry Alignment. 2DGS [16] introduces a normal consistency loss that aligns the normals of Gaussian primitives with those derived from the rendered depth map, ensuring that each 2D splat locally approximates the underlying object surface. However, in boundary regions, the Gaussian primitives often exhibit ambiguous normal directions due to insufficient local support, which can lead to inaccurate geometry reconstruction across different surfaces. To address this issue, we utilize image edges as proxies for geometric edges, assuming that areas with strong image gradients are likely to correspond to surface discontinuities. Thus, we adopt an edge-aware normal consistency loss defined as:
@@
where $\delta=(1-\nabla\bm{I})^{2}$ serves as a per-pixel weight [4] that downweights loss contributions from edge regions, and $\mathcal{I}$ denotes the set of image pixels. $\tilde{\bm{N}}$ is the rendered normal, and $\hat{\bm{N}}$ is the normal estimated from the depth map gradient [16]. To compute normal $\hat{\bm{N}}$, we first project four neighboring depth samples into 3D points in the camera coordinate system. We then estimate the surface normal at pixel $\bm{p}$ by computing the cross product of vectors formed from these projected points, effectively fitting a local plane.
@@ 4.2 Multi-View Alignment
Multi-View Photometric Alignment. While image reconstruction and geometry alignment losses help reduce artifacts and preserve coarse geometry, they often fail to capture fine details. To address this, we draw inspiration from traditional multi-view stereo (MVS) methods [37, 3, 12], which refine surfaces by enforcing photometric consistency across views. Specifically, they project 3D points derived from depth maps onto multiple views and compare their colors to evaluate consistency. By introducing a photometric consistency loss based on plane patches, we leverage multi-view observations to resolve geometric ambiguities, particularly at object boundaries, and enhance reconstruction accuracy.
@@
where $\mathcal{C}(\cdot)$ is the normalized cross-correlation [52], and $V$ is the number of visible pixels. The visibility term $\upsilon_{rs}(\bm{p}_{r})$ indicates whether $\bm{p}_{r}$ is visible in the source view, and $\omega(\bm{p}_{r})$ is a weight accounting for geometric occlusion. Note that we aggregate the losses from all source views by summation, not averaging. The definitions of $\upsilon_{rs}(\bm{p}_{r})$ and $\omega(\bm{p}_{r})$ are detailed in the following.
@@
$\bullet$ During projection via the homography matrix, some pixels may be occluded or exhibit significant geometric error [4]. To avoid the influence of such outliers, we exclude them from the multi-view alignment loss using an occlusion-aware weight. Given a reference 3D point $\bm{x}_{r}$ and its corresponding rendered (or interpolated) depth $\bm{z}_{s}$ in the source view, we first compute the projection error at $\bm{p}_{r}$ as:
@@
Multi-View Feature Alignment. The previously introduced image reconstruction and photometric alignment losses help preserve the shape and structure of the objects. However, image-based losses are susceptible to noise, blur, and low-texture regions. Additionally, due to lighting variations, the color of the same surface point may differ across views, making photometric consistency unreliable. To address these limitations, we introduce a multi-view feature alignment loss. We extract image features using a pretrained network $f$ [58], i.e., $\bm{F}=f(\bm{I})$. Let $\bm{F}_{r}$ denote the reference view’s feature map, and $\bm{F}_{s}$ be one of the source view features, with $\bm{F}_{s}\in\{\bm{F}_{s,i}~|~i=1,2,\ldots,N\}$. Then the pixel-wise feature alignment loss is defined as:
@@
Implicit NeRF [31] 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49 >>12h
VolSDF [48] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 >>12h
NeuS [44] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 >>12h
NeuralWarp [9] 0.49 0.71 0.38 0.38 0.79 0.81 0.82 1.20 1.06 0.68 0.66 0.74 0.41 0.63 0.51 0.68 >>10h
Neuralangelo [27] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 >>12h
PSDF [39] 0.36 0.60 0.35 0.36 0.70 0.61 0.49 1.11 0.89 0.60 0.47 0.57 0.30 0.40 0.37 0.55 -
Explicit 3DGS [21] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 3.4m
SuGaR [14] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 1h
GaussianSurfels[8] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 4.5m
2DGS [16] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 5.8m
GS-Pull [60] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75 5.6m
GOF [56] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 32m
RaDe-GS [57] 0.46 0.73 0.33 0.38 0.79 0.75 0.76 1.19 1.22 0.62 0.70 0.78 0.36 0.68 0.47 0.68 6.5m
PGSR [4] 0.34 0.58 0.29 0.29 0.78 0.58 0.54 1.01 0.73 0.51 0.49 0.69 0.31 0.37 0.38 0.53 15m
GausSurf [43] 0.35 0.55 0.34 0.34 0.77 0.58 0.51 1.10 0.69 0.60 0.43 0.49 0.32 0.40 0.37 0.52 -
@@
Implicit NeuS [44] 0.29 0.29 0.17 0.83 0.24 0.45 0.38 >>12h
Geo-Neus [12] 0.33 0.26 0.12 0.72 0.20 0.45 0.35 >>12h
Neuralangelo [27] 0.70 0.36 0.28 0.89 0.32 0.48 0.50 >>12h
PSDF [39] 0.62 0.39 0.42 0.79 0.47 0.53 0.53 -
Explicit 3DGS [21] 0.13 0.08 0.09 0.04 0.01 0.19 0.09 7.5m
DN-Splatter [42] 0.15 0.11 0.07 0.18 0.01 0.20 0.12 20m
SuGaR [14] 0.14 0.16 0.08 0.33 0.15 0.26 0.19 2h
GaussianSurfels [8] 0.24 0.22 0.07 0.39 0.12 0.24 0.21 5m
2DGS [16] 0.41 0.23 0.16 0.51 0.17 0.45 0.32 7.5m
GS-Pull [60] 0.60 0.37 0.16 0.71 0.22 0.52 0.43 18m
GOF [56] 0.51 0.41 0.28 0.68 0.28 0.59 0.46 40m
RaDe-GS [57] 0.43 0.32 0.21 0.69 0.25 0.51 0.40 9m
PGSR [4] 0.66 0.44 0.20 0.81 0.33 0.66 0.52 25.5m
GausSurf [43] 0.50 0.42 0.30 0.73 0.39 0.65 0.50 -
@@
NeRF [31] 21.46 0.458 0.515 26.84 0.790 0.370 23.85 0.606 0.451
Deep Blending [15] 21.54 0.524 0.364 26.40 0.844 0.261 23.70 0.666 0.318
Instant NGP [32] 22.90 0.566 0.371 29.15 0.880 0.216 25.68 0.706 0.302
MERF [35] 23.19 0.616 0.343 27.80 0.855 0.271 25.24 0.722 0.311
BakedSDF [49] 22.47 0.585 0.349 27.06 0.836 0.258 24.51 0.697 0.309
Mip-NeRF 360 [2] 24.47 0.691 0.283 31.72 0.917 0.180 27.69 0.791 0.237
3DGS [21] 24.64 0.731 0.234 30.41 0.920 0.189 27.20 0.815 0.214
SuGaR [14] 22.93 0.629 0.356 29.43 0.906 0.225 25.82 0.752 0.298
2DGS [16] 24.34 0.717 0.246 30.40 0.916 0.195 27.03 0.805 0.223
GS-Pull [60] 23.76 0.703 0.278 30.78 0.925 0.182 26.88 0.802 0.235
GOF [56] 24.82 0.750 0.202 30.79 0.924 0.184 27.47 0.827 0.194
RaDe-GS [57] 25.17 0.764 0.199 30.74 0.928 0.165 27.65 0.837 0.184
PGSR [4] 24.45 0.730 0.224 30.41 0.930 0.161 27.10 0.819 0.196
GausSurf [43] 25.09 0.753 0.212 30.05 0.920 0.183 27.29 0.827 0.199

### v2

@@ 1 Introduction
Existing methods have attempted to enhance the geometric capabilities of Gaussian splatting. For example, SuGaR [guedon2024sugar] constructs a density field from Gaussians and extracts meshes via level-set searching, but it struggles with large smooth surfaces and is computationally expensive. 2DGS [huang20242d] models scenes using 2D oriented planar Gaussian disks, which inherently represent surfaces and provide view-consistent geometry. However, 2DGS has difficulty reconstructing background geometry and often produces incomplete or distorted surfaces in complex or unbounded scenes. GOF [yu2024gaussian] constructs an opacity field from Gaussians and extracts surfaces using Marching Tetrahedra [doi1991efficient], yielding adaptive mesh resolution without volumetric fusion. Nonetheless, thin structures can be lost and strong lighting contrasts still cause artifacts. GS-Pull [zhang2024neural] integrates a neural signed distance field (SDF), dynamically pulling Gaussians toward the zero-level set of the learned SDF. While this improves surface completeness, it introduces additional network complexity, produces overly smooth surfaces, and primarily focuses on foreground object reconstruction. PGSR [chen2024pgsr] fits Gaussians to local planar hypotheses and uses unbiased depth rendering to improve geometric accuracy. However, it does not fully resolve the challenges posed by complex lighting and remains sensitive to boundary ambiguities in non-planar regions. Overall, previous methods have introduced geometric regularizers or hybrid representations and achieved significant progress. However, they still struggle to address two persistent challenges: illumination-induced artifacts (e.g., shadows and specular highlights) and accurate surface boundary delineation, as shown in Fig. 1. Illumination effects distort photometric losses, while ambiguous boundaries often result in geometry drift or holes.
@@ 2 Related Work
View Synthesis and Gaussian Splatting. Neural Radiance Fields (NeRF) [mildenhall2020nerf] pioneered high-fidelity novel view synthesis by representing a scene as a continuous volumetric density and view-dependent color field, optimized via differentiable volume rendering. Subsequent works accelerated training and rendering through hybrid representations such as multi-resolution hash grids [muller2022instant], explicit voxel or sparse tensor grids [sun2022direct, barron2022mip], and learned feature planes [yu2021plenoctrees]. However, these volumetric methods still entail high memory and computational costs. 3DGS [kerbl20233d] departs from dense volumes by modeling a scene as a sparse cloud of anisotropic 3D Gaussians. Follow-up work has enhanced visual fidelity through anti-aliasing and level-of-detail control [yu2024mip, song2024sa], improved training speed and robustness under sparse views using density regularization and learned radiance priors [yang2024spectrally, niemeyer2024radsplat], and extended 3DGS to dynamic scenes [luiten2023dynamic, yang2023gs4d], relighting [gao2023relightable], and animation [ye2023animatable]. Geometry-aware variants such as FatesGS [huang2025fatesgs], DNGaussian [li2024dngaussian] and GeoGaussian [li2024geogaussian] address sparse-view and textureless regions, while methods like Instantsplat [fan2024instantsplat] and Scaffold-GS [lu2024scaffold] accelerate convergence by leveraging pretraining or hybrid implicit-explicit designs. Despite these advances, most 3DGS variants primarily emphasize appearance quality and lack mechanisms to enforce explicit surface geometry, motivating dedicated reconstruction techniques.
@@
Surface Reconstruction with Gaussians. Extracting accurate surfaces from a 3DGS representation is challenging due to its unstructured nature and supervision based solely on RGB signals. Early approaches convert Gaussians into volumetric density or opacity fields: SuGaR [guedon2024sugar] builds a density field and applies level-set search with Poisson reconstruction [kazhdan2006poisson], but it struggles to recover large, smooth surfaces; GOF [yu2024gaussian] accumulates per-view alpha values into an opacity volume and extracts iso-surfaces with Marching Tetrahedra [doi1991efficient], achieving adaptive resolution but often missing fine, thin structures under high lighting contrast. Other methods project Gaussians into oriented 2D disks (surfels) and fuse via Truncated Signed Distance Function (TSDF) fusion [newcombe2011kinectfusion] or Poisson reconstruction. 2DGS [huang20242d] and GSurfels [dai2024high] improve local alignment but tend to introduce distortions in unbounded scenes and result in incomplete background geometry. PGSR [chen2024pgsr] fits Gaussians to planar patches and adds multi-view photometric and geometric regularization, excelling on planar man-made scenes but remaining sensitive to non-planar boundaries. More recent works [yu2024gsdf, zhang2024neural, lyu20243dgsr, baixin2024gsurf, Li2024MonoGSDFEM, li2025gaussianudf] integrate Signed Distance Fields (SDF) to guide Gaussian placement. GSDF [yu2024gsdf] and 3DGSR [lyu20243dgsr] jointly optimize a neural SDF branch alongside Gaussian parameters using volume-rendered depth and normal supervision, which improves surface smoothness but requires additional network branches. GS-Pull [zhang2024neural] leverages SDF gradients to pull Gaussians toward the zero-level set, enhancing alignment at the cost of limiting object-level reconstruction and producing overly smooth results. Methods that incorporate depth or normal estimators [chen2024vcr, zhang2024rade, turkulainen2024dn, wolf2024surface, wang2024gaussurf] impose priors on Gaussians but rely on TSDF fusion’s fixed resolution or Poisson reconstruction’s sensitivity to noisy inputs, and often struggle under varying illumination or around complex geometric boundaries. Our approach enforces view consistency through multi-faceted constraints during Gaussian optimization, enabling high-fidelity mesh extraction even under challenging lighting and boundary conditions.
@@ 3 Preliminaries
3DGS [kerbl20233d] explicitly represents a scene as a collection of anisotropic 3D Gaussians, which can be rendered to images from arbitrary viewpoints using a splatting-based rasterization technique [zwicker2001ewa]. Specifically, each 3D Gaussian $\mathbb{G}$ is defined as:
@@
Normal and Depth Estimation from Gaussians. The covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ of a 3D Gaussian can be decomposed into a rotation matrix $\bm{R}$ and a scaling matrix $\bm{S}$, i.e., $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$, where $\bm{R}$ contains the three orthogonal eigenvectors, and $\bm{S}$ encodes the scale along these directions. This decomposition resembles an ellipsoid representation: the eigenvectors define the axes of the ellipsoid, while the scale values correspond to the axis lengths. As optimization progresses, the initially spherical Gaussian flattens and approaches a plane [jiang2023gaussianshader]. We take the direction corresponding to the smallest scale factor as the normal $\bm{n}$ of the Gaussian. The distance from the local plane to the camera center is then computed as $\bm{d}=(\bm{R}_{c}^{\top}(\bm{\mu}-\bm{T}_{c}))^{\top}(\bm{R}_{c}^{\top}\bm{n})$, where $\bm{R}_{c}$ is the rotation from the camera to the world frame, and $\bm{T}_{c}$ is the camera center in world coordinates. Given the normal and distance, the depth is obtained by intersecting the viewing ray with the local plane: $\bm{z}=\bm{d}/(\bm{R}_{c}^{\top}\bm{n}~\bm{K}^{-1}\bar{\bm{p}})$, where $\bar{\bm{p}}$ is the homogeneous coordinate of the pixel (we use $\bm{p}$ to denote both the homogeneous and 2D pixel coordinates for simplicity), and $\bm{K}$ is the intrinsic matrix of the camera. Finally, the per-pixel distance, depth, and normal maps under the current viewpoint are rendered using $\alpha$-blending as defined in Eq. (2), where the attribute color $\bm{c}_{i}$ is replaced with the corresponding Gaussian attributes.
@@ 4.1 Single-View Alignment
Edge-aware Image Reconstruction. The original 3DGS [kerbl20233d] and its variants typically employ a color rendering loss, which combines the L1 reconstruction error with a D-SSIM term. While effective for overall image quality, this loss alone is insufficient for accurately capturing object boundaries during surface reconstruction, and it tends to overly smooth high-frequency regions and complex structures. To address this limitation, we propose an edge-aware image reconstruction loss that encourages the model to better preserve sharp structures and boundary details:
@@
Normal-based Geometry Alignment. 2DGS [huang20242d] introduces a normal consistency loss that aligns the normals of Gaussian primitives with those derived from the rendered depth map, ensuring that each 2D splat locally approximates the underlying object surface. However, in boundary regions, the Gaussian primitives often exhibit ambiguous normal directions due to insufficient local support, which can lead to inaccurate geometry reconstruction across different surfaces. To address this issue, we utilize image edges as proxies for geometric edges, assuming that areas with strong image gradients are likely to correspond to surface discontinuities. Thus, we adopt an edge-aware normal consistency loss defined as:
@@
where $\delta=(1-\nabla\bm{I})^{2}$ serves as a per-pixel weight [chen2024pgsr] that downweights loss contributions from edge regions, and $\mathcal{I}$ denotes the set of image pixels. $\tilde{\bm{N}}$ is the rendered normal, and $\hat{\bm{N}}$ is the normal estimated from the depth map gradient [huang20242d]. To compute normal $\hat{\bm{N}}$, we first project four neighboring depth samples into 3D points in the camera coordinate system. We then estimate the surface normal at pixel $\bm{p}$ by computing the cross product of vectors formed from these projected points, effectively fitting a local plane.
@@ 4.2 Multi-View Alignment
Multi-View Photometric Alignment. While image reconstruction and geometry alignment losses help reduce artifacts and preserve coarse geometry, they often fail to capture fine details. To address this, we draw inspiration from traditional multi-view stereo (MVS) methods [schonberger2016structure, campbell2008using, fu2022geo], which refine surfaces by enforcing photometric consistency across views. Specifically, they project 3D points derived from depth maps onto multiple views and compare their colors to evaluate consistency. By introducing a photometric consistency loss based on plane patches, we leverage multi-view observations to resolve geometric ambiguities, particularly at object boundaries, and enhance reconstruction accuracy.
@@
where $\mathcal{C}(\cdot)$ is the normalized cross-correlation [yoo2009fast], and $V$ is the number of visible pixels. The visibility term $\upsilon_{rs}(\bm{p}_{r})$ indicates whether $\bm{p}_{r}$ is visible in the source view, and $\omega(\bm{p}_{r})$ is a weight accounting for geometric occlusion. Note that we aggregate the losses from all source views by summation, not averaging. The definitions of $\upsilon_{rs}(\bm{p}_{r})$ and $\omega(\bm{p}_{r})$ are detailed in the following.
@@
$\bullet$ During projection via the homography matrix, some pixels may be occluded or exhibit significant geometric error [chen2024pgsr]. To avoid the influence of such outliers, we exclude them from the multi-view alignment loss using an occlusion-aware weight. Given a reference 3D point $\bm{x}_{r}$ and its corresponding rendered (or interpolated) depth $\bm{z}_{s}$ in the source view, we first compute the projection error at $\bm{p}_{r}$ as:
@@
Multi-View Feature Alignment. The previously introduced image reconstruction and photometric alignment losses help preserve the shape and structure of the objects. However, image-based losses are susceptible to noise, blur, and low-texture regions. Additionally, due to lighting variations, the color of the same surface point may differ across views, making photometric consistency unreliable. To address these limitations, we introduce a multi-view feature alignment loss. We extract image features using a pretrained network $f$ [zhang2023vis], i.e., $\bm{F}=f(\bm{I})$. Let $\bm{F}_{r}$ denote the reference view’s feature map, and $\bm{F}_{s}$ be one of the source view features, with $\bm{F}_{s}\in\{\bm{F}_{s,i}~|~i=1,2,\ldots,N\}$. Then the pixel-wise feature alignment loss is defined as:
@@
Implicit NeRF [mildenhall2020nerf] 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49 >>12h
VolSDF [yariv2021volume] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 >>12h
NeuS [wang2021neus] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 >>12h
NeuralWarp [darmon2022improving] 0.49 0.71 0.38 0.38 0.79 0.81 0.82 1.20 1.06 0.68 0.66 0.74 0.41 0.63 0.51 0.68 >>10h
Neuralangelo [li2023neuralangelo] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 >>12h
PSDF [su2024psdf] 0.36 0.60 0.35 0.36 0.70 0.61 0.49 1.11 0.89 0.60 0.47 0.57 0.30 0.40 0.37 0.55 -
Explicit 3DGS [kerbl20233d] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 3.4m
SuGaR [guedon2024sugar] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 1h
GaussianSurfels[dai2024high] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 4.5m
2DGS [huang20242d] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 5.8m
GS-Pull [zhang2024neural] 0.51 0.56 0.46 0.39 0.82 0.67 0.85 1.37 1.25 0.73 0.54 1.39 0.35 0.88 0.42 0.75 5.6m
GOF [yu2024gaussian] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 32m
RaDe-GS [zhang2024rade] 0.46 0.73 0.33 0.38 0.79 0.75 0.76 1.19 1.22 0.62 0.70 0.78 0.36 0.68 0.47 0.68 6.5m
PGSR [chen2024pgsr] 0.34 0.58 0.29 0.29 0.78 0.58 0.54 1.01 0.73 0.51 0.49 0.69 0.31 0.37 0.38 0.53 15m
GausSurf [wang2024gaussurf] 0.35 0.55 0.34 0.34 0.77 0.58 0.51 1.10 0.69 0.60 0.43 0.49 0.32 0.40 0.37 0.52 -
@@
Implicit NeuS [wang2021neus] 0.29 0.29 0.17 0.83 0.24 0.45 0.38 >>12h
Geo-Neus [fu2022geo] 0.33 0.26 0.12 0.72 0.20 0.45 0.35 >>12h
Neuralangelo [li2023neuralangelo] 0.70 0.36 0.28 0.89 0.32 0.48 0.50 >>12h
PSDF [su2024psdf] 0.62 0.39 0.42 0.79 0.47 0.53 0.53 -
Explicit 3DGS [kerbl20233d] 0.13 0.08 0.09 0.04 0.01 0.19 0.09 7.5m
DN-Splatter [turkulainen2025dn] 0.15 0.11 0.07 0.18 0.01 0.20 0.12 20m
SuGaR [guedon2024sugar] 0.14 0.16 0.08 0.33 0.15 0.26 0.19 2h
GaussianSurfels [dai2024high] 0.24 0.22 0.07 0.39 0.12 0.24 0.21 5m
2DGS [huang20242d] 0.41 0.23 0.16 0.51 0.17 0.45 0.32 7.5m
GS-Pull [zhang2024neural] 0.60 0.37 0.16 0.71 0.22 0.52 0.43 18m
GOF [yu2024gaussian] 0.51 0.41 0.28 0.68 0.28 0.59 0.46 40m
RaDe-GS [zhang2024rade] 0.43 0.32 0.21 0.69 0.25 0.51 0.40 9m
PGSR [chen2024pgsr] 0.66 0.44 0.20 0.81 0.33 0.66 0.52 25.5m
GausSurf [wang2024gaussurf] 0.50 0.42 0.30 0.73 0.39 0.65 0.50 -
@@
NeRF [mildenhall2020nerf] 21.46 0.458 0.515 26.84 0.790 0.370 23.85 0.606 0.451
Deep Blending [hedman2018deep] 21.54 0.524 0.364 26.40 0.844 0.261 23.70 0.666 0.318
Instant NGP [muller2022instant] 22.90 0.566 0.371 29.15 0.880 0.216 25.68 0.706 0.302
MERF [reiser2023merf] 23.19 0.616 0.343 27.80 0.855 0.271 25.24 0.722 0.311
BakedSDF [yariv2023bakedsdf] 22.47 0.585 0.349 27.06 0.836 0.258 24.51 0.697 0.309
Mip-NeRF 360 [barron2022mip] 24.47 0.691 0.283 31.72 0.917 0.180 27.69 0.791 0.237
3DGS [kerbl20233d] 24.64 0.731 0.234 30.41 0.920 0.189 27.20 0.815 0.214
SuGaR [guedon2024sugar] 22.93 0.629 0.356 29.43 0.906 0.225 25.82 0.752 0.298
2DGS [huang20242d] 24.34 0.717 0.246 30.40 0.916 0.195 27.03 0.805 0.223
GS-Pull [zhang2024neural] 23.76 0.703 0.278 30.78 0.925 0.182 26.88 0.802 0.235
GOF [yu2024gaussian] 24.82 0.750 0.202 30.79 0.924 0.184 27.47 0.827 0.194
RaDe-GS [zhang2024rade] 25.17 0.764 0.199 30.74 0.928 0.165 27.65 0.837 0.184
PGSR [chen2024pgsr] 24.45 0.730 0.224 30.41 0.930 0.161 27.10 0.819 0.196
GausSurf [wang2024gaussurf] 25.09 0.753 0.212 30.05 0.920 0.183 27.29 0.827 0.199